Wednesday, 26 August

Wednesday, 26 August

Short Communications and Poster Sessions

Short Communications: Wednesday, 26 August, 14:00–14:45

14:00–14:15  14:15–14:30  14:30–14:45

Sect.: 03

Function Fields and Curves II, Chair: Dmitry Logachev

MA 311 Sudhir R. Ghorpade Number of Points of Singular Varieties over Finite Fields (p.37)

Seman Wong

On the density of elliptic curves (p.47)

Dmitry Logachev

Mazur–Bloch method for Siegel threefolds (p.41)

Sect.: 04

Surfaces/Higher-Dimensional Varieties IV, Chair: Viktor S. Kulikov

MA 042 Yuri G. Prokhorov Mori contractions of relative dimension 1

(p.59)

Maria Cristina Ronconi

Algebraic threefolds with q_{1} = q_{2} = p_{g} = 0, P_{2i} = 1, and P_{2i+1} = 0 (p.60)

Ezio Stagnaro

Pluricanonical maps of two general type threefolds with p_{g} = 0 (p.60)

Sect.: 06

Algebraic Topology of Manifolds, Chair: Lawrence W. Conlon

H 112 Ulrich Koschorke μ-invariants of link maps in arbitrary manifolds (p.90)

Yuri V. Muranov

Surgery and splitting obstruction groups for one-sided submanifolds (p.94)

Gerard A. Venema

Representing homology classes of 4manifolds (p.99)

Sect.: 08

Special Functions, Chair: Promarz M. Tamrazov

MA 041 Raoul R.F.G. Gloden Propriétés des polynômes orthogonaux. Développements de cas particuliers

(p.126)

Kathy A. Driver

Zeros of Hypergeometric Functions (p.121)

Momčilo Bjelica

Asymptotic of means (p.117)

Sect.: 08

Metrizable Spaces, Riesz Spaces, and Integral Representations, Chair: Paul Ressel

MA 309 Shozo Koshi Partially ordered linear space (p.136)

Luis M. Sánchez Ruiz

On Metrizable Barreled Spaces, Cardinals and Dimensionality (p.153)

Paul Ressel

Semigroup representations, positive definite functions, and abelian C*-algebras

(p.152)

Sect.: 08

Operator Algebras and Systems, Chair: Palle E.T. Jorgensen

MA 043 B.V. Rajarama Bhat Minimal dilation of quantum dynamical semigroups to semigroups of endomorphisms (p.116)

Wend Werner

Operator Systems without Unit (p.163)

Palle E.T. Jorgensen

Ruelle operators: Functions which are harmonic with respect to a transfer operator

(p.133)

Sect.: 10

PDE and Dynamical Systems, Chair: Alexandre Kometch

H 111 Andrey Yu. Goritsky Integral manifolds with exponential tracking (p.204)

Jörg Härterich

Heteroclinic Orbits Between Rotating Waves in Hyperbolic Balance Laws

(p.205)

Alexandre Kometch

Attractors for Hamiltonian Nonlinear Wave Equations (p.210)

Sect.: 10

Non-linear Theory, Chair: Masafumi Yoshino

H 110 Vladimir Kozlov Singularities of solutions of nonlinear elliptic equations (p.211)

Sidi Mohammed Bouguima

Free boundary problem in the plane (p.199)

Masafumi Yoshino

Global solvability of fully nonlinear equations and Toeplitz operators (p.225)

Sect.: 12

Gaussian Processes, Chair: Manfred Breger

H 107 Krzysztof Dȩbicki Asymptotic of the tail distribution of supremum from a Gaussian process (p.250)

Tudor Sireteanu

An extension of Gaussian equivalent linearization (p.270)

Victoria Parkhomenko

Spectral representation of the scaling limit solutions of the Burgers’ equation with random data (p.265)

Sect.: 12

Nonparametric Curve Estimation, Chair: Joerg Polzehl

H 1029 Renfang Jiang The mean sum of squared error of a wavelet based multinomial density estimator (p.255)

Eva Herrmann

Visualization of empirical test results for pseudorandom number generators (p.253)

Hua Liang

Partially linear errors-in-variables model (p.258)

Sect.: 16

Optimization, Chair: Uwe T. Zimmermann

H 106 Ludwig Cromme Optimization of a Production Line in Brown Coal Industry (p.321)

Ludwig Kohaupt

Optimization for an Exhaust Emission Problem by the Simplex Method — A Report on Work from the Automotive Industry - (p.326)

Uwe T. Zimmermann

Hard Integer Programming Problems from Railroad Planning (p.334)

Sect.: 17

Stochastic Programming II, Chair: Vladimir Norkin

H 2033 René Henrion Quantitative stability of convex stochastic programs with probabilistic constraints

(p.342)

Werner Römisch

Perturbation Analysis of Stochastic Programming Problems (p.349)

Roger Fandom Noubiap

A Minimax Algorithm for Calculating Statistical Tests under Generalized Moment Conditions (p.340)

Sect.: 17

Miscellaneous I, Chair: Mohamed Ali Hammami

H 2013 Tomáš Roubíček Relaxation in optimization theory and variational calculus Mohamed Ali Hammami Stabilization of a certain class of nonlinear systems via an observer

Short Communications: Wednesday, 26 August, 15:00–15:45

15:00–15:15  15:15–15:30  15:30–15:45

Sect.: 02

Algebras III, Chair: Shavkat A. Ayupov

MA 311 Bernhard Drabant Cross Product Bialgebras (p.17)

Alexander A. Mikhalev

Free Leibniz algebras (p.23)

Daouda Sangare

Hilbert quasi-polynomial functions and generalized analytic spread (p.27)

Sect.: 02

Categories, Chair: Reinhard Börger

MA 042 Hvedri Inassaridze Karoubi’s Conjecture for polynomial extensions of stable C*-algebras (p.19)

Fred E.J. Linton

Banach Spaces Monadic over Free Pointed Cones (p.22)

Siaka Kante Bamba

On quadratic modules (p.13)

Sect.: 06

Lusternik-Shnirelmann Category, Chair: Hans Scheerer

H 112 Jean-Michel Lemaire Rational Lusternik-Schnirelmann category of Poincaré complexes (p.91)

Yuli B. Rudyak

Lusternik–Schnirelmann category of symplectic manifolds and the Arnold conjecture (p.96)

Sect.: 08

Kleinian Groups, Chair: Alexander Jur’evich Vasil’ev

MA 041 Ruben A. Hidalgo A_{4}, A_{5}, and S_{4} of Schottky Type (p.130)

Weiqi Yang

Discrete groups and dense groups (p.164)

Javier Pérez

Harmonic differentials on Klein Surfaces (p.147)

Sect.: 08

Analytic Functions and Toeplitz Operators, Chair: Marek Ptak

MA 309 Ali Abkar A comparison theorem concerning hyperinvariant subspaces on finitely connected domains (p.113)

Slavko Simic

Entire Functions of Finite Order and Regularly Varying Sequences (p.156)

Marek Ptak

Toeplitz operators; dichotomy of reflexivity and transitivity (p.149)

Sect.: 08

Operator Algebras and Algebras of Operators, Chair: Vladimir Manuilov

MA 043 Raphael S. Ponge Wodzicki Residue for Heisenberg Manifolds (p.148)

Georgios F. Nassopoulos

Coproducts and projective presentations of operator algebras (p.144)

Vladimir Manuilov

Almost representations of discrete groups and K-theory (p.140)

Sect.: 09

Nonlinear Evolution Equations, Chair: Michael V. Pomazanov

H 1029 Valery A. Gaiko Global bifurcation theory and multiple limit cycles Samuel S. Okoya A criterion for solutions of a model quasilinear parabolic equation Michael V. Pomazanov About the structure of some typical singularities in implicit ordinary differential equations

Sect.: 10

Hyperbolic Equations IV (General Theory), Chair: Pavel A. Krutitskii

H 111 Ludmila S. Pulkina A nonlocal problem for the hyperbolic equation (p.217)

Peetta Kandy Ratnakumar

Wave Equation for the Hermite Operator (p.218)

Pavel A. Krutitskii

Initial-boundary value problems in multiply connected domains for an equation of composite type (p.212)

Sect.: 10

Applied Problems II (Elasticity & Fusion), Chair: Juan Francisco Padial

H 110 Alexander A. Kosmodemyanskii On the location of fail points in Torsion Problem (p.211)

Juan Francisco Padial

Some mathematical models related to nuclear fusion (p.216)

Sect.: 11

Algebraic Methods in Physics, Chair: Victor V. Zharinov

H 107 Andrej V. Bulinskij On dynamical semigroups compatible with states of von Neumann algebras (p.229)

Fernando F. Lledó

On some algebraic properties of massless free nets (p.239)

Victor V. Zharinov

On derivations of the Heisenberg algebra (p.246)

Sect.: 16

Numerical Applications of Matrix Theory, Chair: Dmitrii O. Logofet

H 106 Igor N. Lyashenko � ��������������� ��� ����� ��� ���������˘� ������ ���������- ����� (p.329)

Dmitrii O. Logofet

When matrix “flowers” help proving theorems on stability (p.328)

Gennady Victorovich Alferov

Motion Planning for Multi — Agent Systems in the Time — Varying Environments

Sect.: 17

Semi-infinite and Semi-definite Optimization, Chair: René Henrion

H 2033 Jan-J. Ruckmann On Generalized Semi-infinite Optimization Problems (p.349)

Oliver Stein

On level sets of marginal functions (p.351)

Ivo Nowak

Bézier-Semidefinite Relaxations of nonconvex quadratic programs over a simplex

(p.346)

Sect.: 17

Miscellaneous II, Chair: Roland Wessäly

H 2013 Ingo Bremer Experiences with the Optimization of Batch Distillation Processes Rafail Gabasov Synthesis of optimal dynamical systems Roland Wessäly New inequalities for a capacitated survivable network design problem (p.355)

Short Communications: Wednesday, 26 August, 16:15–17:00

16:15–16:30  16:30–16:45  16:45–17:00

Sect.: 02

Linear Algebra and Quadratic Forms, Chair: Stefan Heiss

MA 311 Herbert Möller New central algorithms in linear algebra

(p.23)

Dilip N. Sheth

Completion of a unimodular vector to an invertible matrix (p.28)

Ulf Rehmann

Higher anisotropic kernels of quadratic forms (p.26)

Sect.: 04

Algebraic Cycles, Chair: Yuri G. Prokhorov

MA 042 Manfred F.G. Streit Field of Definition and Galois Orbits of some Non–Congruence Subgroups (p.61)

Serguei G. Tankeev

Algebraic cycles of small codimension on simple 2p- or 4p-dimensional abelian variety (p.62)

Johannes Huisman

Uniformization of real algebraic curves and complex analytic coordinates on Teichmüller spaces (p.54)

Sect.: 05

Analysis on Manifolds, Chair: Marian Gidea

MA 041 Yuri G. Borisovich Methods of global analysis in problems of solvability of nonlinear equations and inclusions (p.65)

Sylvie Paycha

Regularization as a Geometric and Stochastic Tool in Infinite Dimensions

(p.78)

Marian Gidea

Non-Smooth Dynamical Systems that Exhibit Hyperbolic Behavior (p.69)

Sect.: 05

Polyhedra/Webs, Chair: Hakki Ismail Erdogan

MA 043 Anatoly D. Milka Linear isometric deformations of polyhedra

(p.77)

Hakki Ismail Erdogan

On a class of hexagonal-surface-web of spheres (p.68)

Sect.: 09

Complex Dynamical Systems, Chair: John C. Mayer

H 111 Toshiyuki Sugawa An explicit bound for uniform perfectness of Julia sets (p.193)

Hiroki Sumi

Dynamics of rational semigroups and Hausdorff dimension of the Julia sets

(p.193)

John C. Mayer

Invariant subcontinua in Siegel and Cremer Julia sets of polynomials (p.183)

Sect.: 09

Impulsive Systems, Chair: Anatoli Myshkis

H 110 Dariusz Idczak Stability in semilinear problems (p.177)

Cristian Dascalu

Antiplane Shear Instability in Elastodynamics with Slip Dependent Friction

(p.174)

Anatoli Myshkis

Autonomous impulsive-differential equations (p.184)

Sect.: 11

Field Theory III, Chair: Henrik Petersson

H 107 Robert Fittler Nonstandard Quantumelectrodynamics

(p.233)

Takashi Kimura

Cohomological Field Theories and Tautological Classes (p.236)

Henrik Petersson

Properties of analytical maps on infinite dimensional locally convex spaces (p.242)

Sect.: 13

Combinatorial Constructions, Chair: Mark Kayll

H 106 Abdelkader Khelladi Some Results on the Fiber Product of Graphs (p.278)

Andrew J. Woldar

General properties of algebraically defined graphs (p.287)

Mark Kayll

Stochastic independence properties of hard-core distributions (p.278)

Sect.: 15

Wavelets IV (organized by Andreas Rathsfeld), Chair: Andreas Rathsfeld

H 2033 Jürgen Prestin Polynomial wavelets Andreas Rathsfeld Piecewise linear wavelet collocation for boundary element methods (p.310)

Sect.: 19

Miscellaneous, Chair: Seemanapalli V. Sastry

H 2013 Seemanapalli V. Sastry History of Mathematics: Yesterday, Today, and Tomorrow (p.368)

Short Communications: Wednesday, 26 August, 17:15–18:00

17:15–17:30  17:30–17:45  17:45–18:00

Sect.: 02

Groups IV, Chair: Bettina Wilkens

MA 311 Alexei N. Krasilnikov Identities in metanilpotent groups (p.21)

Ernest B. Vinberg

On a generalization of Coxeter groups Munibur Rahman Chowdhury A commutativity theorem for groups

(p.15)

Sect.: 05

Differential Equations on Manifolds, Chair: Congming Li

MA 041 Atsushi Tachikawa A Nonexistence Result of Harmonic Maps between Hadamard Manifolds (p.82)

Asya Vasilyevna Aminova

The pseudo-Riemannian manifolds with common geodesics (p.64)

Congming Li

Prescribing Scalar Curvature on S^{n} (p.74)

Sect.: 09

Invariant Manifolds and Normal Forms, Chair: Pei Yu

H 111 Alexander Ya. Kopanskii Normal forms of vector fields satisfying certain geometric conditions (p.180)

Alexander M. Kovalev

The equations of invariant and oriented manifolds of dynamical systems (p.180)

Pei Yu

Simplest Normal Forms of Differential Equations (p.196)

Sect.: 09

Boundary Value Problems, Chair: Guiseppe Anichini

H 110 Nusrat Rajabov Linear conjugate boundary value problems for the second order linear ordinary differential equation with super–singular coefficients (p.186)

Masashi Kisaka

On the Connnected Components of Julia Sets of Transcendental Entire Functions

(p.180)

Guiseppe Anichini

Boundary value problems for implicit ODE’s in a singular case (p.170)

Sect.: 11

Thermodynamics, Hydrodynamics, Non-linear Dynamics IV, Chair: Juan M. Alvaro Vilumbrales

H 107 Sergei Fedotov Travelling Waves in a Reaction-Diffusion System with KPP-kinetics (p.233)

Alexander A Alexeyev

Superposition principles for nonlinear partial differential equations in (1+1) dimension (p.227)

Juan M. Alvaro Vilumbrales

Cylindrical and spherical problems of Stefan: analytical solutions and numerical estimates (p.228)

Sect.: 13

Geometry IV, Chair: Yuri Grigoriyevich Stoyan

H 106 Alexei A. Tuzhilin Planar Local Minimal Networks with Convex Boundaries (p.286)

Alexandr O. Ivanov

Linear Networks in Euclidean Spaces R^{n} (p.277)

Yuri Grigoriyevich Stoyan

Properties of combinatorial set images into the arithmatic Euclidean spaces (p.285)

Poster Session: Wednesday, 26 August, 14:00–15:45

Sect.: 08  Analysis

MA 144 Ants Aasma Summability with speed and comparison of orders of approximation of Fourier expansions by different matrix methods of summability (p.112)

El Hassan E.H. El Kinani

q-Derivation, Sine algebra, and Ut(sl(2)) (p.123)

Stefan B.A. Jakobsson

Applications of weighted Hele-Shaw flow and kernel function techniques to Green functions for weighted biharmonic operators (p.132)

Volodymyr M. Kolodyazhniy Nonlinear Identification Problem Solution Using the Wavelets of the Atomic Functions (p.136)

Hrvoje Kraljevic

Landau’s type inequalities for contraction cosine functions (p.137)

Carlos Lizama

A Characterization of Uniform Continuity for Volterra Equations in Hilbert Spaces

(p.140)

Igor-Ya. Novikov Time-frequency localization of modified Daubechies wavelets (p.144)

Alan L. Schwartz

Hypergroups and their maximum subgroup (p.154)

Katsunori Shimomura

On the transformations which preserve polytemperatures (p.156)

Feng-Yu Wang Harnack inequality for heat semigroups and log-Sobolev inequalities (p.163)

Oxana A. Ziza

On absolute summability of orthogonal series (p.167)

Sect.: 19

History of Mathematics

MA 144 Carlos Gomez Bermudez Le système des nombres chez G. Cantor

(p.366)

Harald Gropp

Levi, Steinitz, Zacharias, and configurations –Three German mathematicians in the 20^{th} century and a combinatorial structure (p.366)

Aleksandar T. Lipkovski

Rational curves with elliptic arcs — Serret’s solution of Legendre’s problem

(p.368)

Man-Keung Siu On the myth of an ancient Chinese theorem about primality (p.369)

Poster Session: Wednesday, 26 August, 16:15–18:00

Sect.: 17

Control Theory and Optimization

MA 144 Shamsha A. Aipanov The solutions of some problems of optimal control of spacecraft (p.335)

Vladimir Borisov

Bundles with chattering extremals (p.337)

Anna Debinska-Nagorska

A Non-linear Hyperbolic Control Problem with Neumann Boundary Condition — Convergence of Galerkin Approximation

(p.338)

Roland Hildebrand Optimal synthesis of controlled systems near a focus (p.343)

Christian Schulz

Optimal scheduling of a chemical batch plant (p.350)

Zdzislaw Stempien

On the Optimal Control of Nonlinear Hyperbolic System — Approximation (p.351)

Pierre von Mouche Sur deux problèmes géométriques dans la théorie des jeux (p.354)

Olga Winzer

Complexity of vector spanning trees problems with topological criteria (p.355)

Stanislav T. Zavalishchin

Dynamic impulse systems. Optimization problems in fluid mechanics (p.356)